# FaithUn: Toward Faithful Forgetting in Language Models by Investigating the Interconnectedness of Knowledge

### v1

@@ 3.3 Dataset Summary
Dataset Format. Each instance of the dataset is denoted as a tuple: $d=\langle\mathcal{C}^{i},\mathcal{C}_{p}^{i},\mathcal{C}_{m}^{i},\mathcal{C}_{s}^{i}\rangle$. The FaithUn dataset starts from a core factual triple $(s,r,o)$, which forms the knowledge of the Base QA dataset $\mathcal{C}^{i}$. There are also the Paraphrased QA dataset $\mathcal{C}_{p}^{i}$, based on the same triple, the Multi-hop QA dataset $\mathcal{C}_{m}^{i}$, which extends from the original triple $(s,r,o)$, and the Same-answer QA dataset $\mathcal{C}_{s}^{i}$, which shares the same answers as the Base QA dataset’s questions but has different contexts. Each of these datasets ($\mathcal{C}^{i}$, $\mathcal{C}_{p}^{i}$, $\mathcal{C}_{m}^{i}$, and $\mathcal{C}_{s}^{i}$) is composed of question-answer pairs $(q,a)$, and they also include false answer options to enable evaluation through Multiple-choice QA (MCQA). The details for the MCQA setting are described in Section 3.4. We also describe detailed examples in Table 10. In addition, we summarize the differences our benchmark addresses compared to existing benchmarks (Shi et al., 2024; Tian et al., 2024; Li et al., 2024a; Maini et al., 2024; Jin et al., 2024) in Table 1.
Dataset Statistics. After collecting triples of the Base QA dataset, we filter only triples including matched Multi-hop QA or Same-answer QA samples. Therefore, each QA instance in the Base QA dataset serves as a cluster for evaluating the faithfulness of unlearning methods. Consequently, we collect 664 QA pairs for the Base QA dataset. Each Base QA instance includes three paraphrased questions, for a total of 1,992 paraphrased QA instances in our dataset. FaithUn also include 1,714 instances for multi-hop QA datasets. Furthermore, our dataset includes 4,671 instances for the Same-answer QA dataset. The statistics of the constructed FaithUn datasets are shown in Table 4.
Dataset Quality. We adopt a ChatGPT variant to generate natural language questions, a commonly used and powerful approach, following existing studies (Shi et al., 2024; Jin et al., 2024; Maini et al., 2024). However, to further investigate the quality of the dataset, we conducted a human evaluation for the generated questions. Specifically, we recruited crowd workers fluent in English through the university’s online community and had them evaluate 800 generated natural language questions. The results revealed an error rate of 0%, confirming the reliability of our benchmark.
@@ 3.4 Evaluation Framework
To evaluate the faithfulness of unlearning methods, we first split the forget set $\mathcal{C}_{f}$, the retaining set $\mathcal{C}_{r}$, and the test set $\mathcal{C}_{t}$ from the entire Base QA dataset $\mathcal{C}$. Then, we train a language model to unlearn the forget set while maintaining knowledge of the retaining set. We further evaluate the unlearned model to the test set to assess knowledge retention for unseen data. In addition, we evaluate the unlearned model with the other constructed datasets (i.e., $\mathcal{C}_{p}$, $\mathcal{C}_{m}$, and $\mathcal{C}_{s}$) mapped to the forget and test sets to analyze the aspect of superficial unlearning.
Our unlearning framework consists of two types of input formats: (1) general QA format, and (2) multiple-choice QA (MCQA) format. We use the general QA format for unlearning and the MCQA format for evaluation. The general QA format inputs a question without an additional template, while the MCQA format uses a template that includes instructions and answer options. Suppose we aim to unlearn the knowledge of the question "Who is the mother of Barack Obama?", then we train a language model not to output the correct answer (i.e., "Stanley Ann Dunham") using only the question as an input. However, many users use a language model with various instruction templates, and an unlearned model should be evaluated in a stricter environment considering generalization. Furthermore, evaluating all possible answers to a question is one of the most challenging aspects of QA evaluation. Therefore, we utilize the MCQA form to evaluate an unlearned model. This makes it easier for LLMs to derive knowledge since they are given answer options; thus, it makes unlearning algorithms harder to apply. For this reason, we use the MCQA setting to evaluate unlearned models in more challenging and practical settings.
@@ 3.5 Evaluation Metrics
We propose various metrics to evaluate the basic unlearning performance and the superficial unlearning performance. We use exact match to calculate the score of all metrics. (1) Unlearning Accuracy (UA): We compute accuracy for the forget set $\mathcal{C}_{f}$ to evaluate the basic unlearning performance. (2) Extended Unlearning Accuracy (UA): We compute accuracy for the Paraphrased QA set $\mathcal{C}_{p}$ to evaluate the generalized unlearning performance. (3) Test Accuracy (TA): We compute accuracy for the test set $\mathcal{C}_{t}$ to evaluate whether knowledge of unseen instances is maintained after the unlearning process. (4) Same-answer Test Accuracy (SA): We compute accuracy for the Same-answer QA set $\mathcal{C}_{s}$ to analyze shortcut unlearning. An unlearning algorithm may only superficially degrade the probability of the answer regardless of context. (5) Multi-hop Test Accuracy (MA): We compute accuracy for $\mathcal{C}_{m}$ matched with each instance of $\mathcal{C}_{f}$ and $\mathcal{C}_{t}$ to evaluate whether the interconnected knowledge of instances is effectively unlearned. To derive the aggregated MA score, we first compute the individual accuracies, MAf for all $(q,a)\in\mathcal{C}_{m}$ mapped to $\mathcal{C}_{f}$ and MAt for all $(q,a)\in\mathcal{C}_{m}$ mapped to $\mathcal{C}_{t}$; then, we compute the aggregated score, MA, by averaging the scores, ($100-$MAf) and MAt. Although the number of samples in $\mathcal{C}_{t}$ is generally higher than in $\mathcal{C}_{f}$, we average the scores with equal weight, as we assume that unlearning samples in $\mathcal{C}_{f}$ is important due to significant privacy concerns. (6) Total Score (Score): We aggregate all the evaluation scores by averaging ($100-$UA), TA, SA, and MA, to present the overall performance.
@@ 4 Method: KLUE
Unlearning methods should erase only the knowledge associated with the target knowledge while preserving all other knowledge. In this section, we describe the method, KLUE, that identifies neurons contextually related to the target knowledge and updates only them during the unlearning process.
@@ 4.1.1 Knowledge Quantification
We utilize an attribution method (Shrikumar et al., 2016) to extract the importance of neurons for specific world knowledge from language models. It is usually used to derive the importance of the input features (i.e., pixel, token) for performing a specific task, but Yang et al. (2023) expands the attribution formula to the importance of intermediate neurons in language models. Formally, suppose we have $P_{\theta}(y|x)=\prod_{t=1}^{T}P_{\theta}(y_{t}|x,y_{1},...,y_{t-1})$ that represents a language model. The contribution of an $i$-th neuron to the representation $h$ in a particular layer, in predicting an answer $a$ given a question $q$ using $P_{\theta}$, is defined as follows:
@@
[…]
where $h^{l}$ means $l$-th token representation of $h$, and $\partial P_{\theta}(a|q)/\partial h^{l}_{i}$ is the gradient of $P_{\theta}(a|q)$ with respect to $h^{l}_{i}$. In this study, we use transformer variants for experiments; thus, activation scores and gradients of a specific layer are computed for each input token representation. Therefore, if an input text includes $L$ tokens, we have $L$ attribution scores for each neuron; thus, we aggregate attributions of tokens by using max aggregation to acquire a single neuron knowledge attribution $A^{(q,a)}_{i}(h)$.
@@ 4.1.2 Superficial Knowledge Regularization
Equation 3 computes the knowledge relevance of each neuron for a specific $(q,a)$ pair. However, this equation may include undesirable information that only serves to increase the likelihood of the answer $a$ regardless of the given context. To eliminate undesirable information from the computed attribution, we construct synthetic mismatched QA pairs $(q^{\prime},a)\in\mathcal{C}^{\prime}$, where the answers remain the same as the target answer $a$, while the questions are randomly sampled independently of the answer. Then, we compute the attribution score for each mismatched pair and average them. Since a question and an answer included in mismatched pairs are contextually irrelevant, the computed attribution corresponds to the degree that unconditionally increases the likelihood of the answer regardless of the context (superficial knowledge). Therefore, we can compute the final knowledge attribution, $\mathcal{I}$, containing only contextual knowledge by excluding the information of the mismatched attribution from the basic knowledge attribution as follows:
$S^{(q,a)}_{i}(h)=\sum_{(q^{\prime},a)\in\mathcal{C}^{{}^{\prime}}}{\tilde{A}^{(q^{\prime},a)}_{i}(h)}$, (4)
@@
where $\mathcal{C}^{{}^{\prime}}$ is a set including mismatched question and answer pairs. $N$ is the number of mismatched samples, and $\alpha$ is a hyper-parameter to determine the magnitude of knowledge exclusion. $\tilde{A}$ means a negative value of $A$ is converted to the zero value. Since the negative values of the attribution score are negative contributions to a specific knowledge, it is reasonable to eliminate that unnecessary information. We use $\mathcal{C}^{f}$ and $\mathcal{C}^{r}$ as a pool to sample mismatched questions. Notice that this regularization enhances the quantification of contextual knowledge; thus, it can improve multi-hop reasoning and mitigate shortcut unlearning.
4.2 Unforgotten Sample-localized Unlearning
If we repeatedly unlearn samples that have already been sufficiently unlearned, it leads to overfitting in language models. Therefore, we select only the samples that are not completely forgotten in the unlearning process to preserve the generalization performance. Specifically, in each epoch’s unlearning process, we select and unlearn only questions that satisfy the knowledge memorization criteria (Described in Section 3.1).
4.3 Knowledge Neuron-localized Unlearning
After selecting unforgotten samples, we localize and update only the knowledge neurons corresponding to those selected samples in the language model. Specifically, we first compute gradients of parameters for the selected unforgotten samples. Then, we quantify the knowledge relevance of each neuron by using the equations 3 and 4, and sort neurons of the whole target layers by the knowledge relevance scores; then, we select the top-$n$ knowledge neurons. We finally mask gradients of the parameters for knowledge-irrelevant neurons to exclude them from the unlearning process. Suppose that a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times k}$ is a linear matrix multiplication parameter of a language model, and the gradient computed for the parameter is $\nabla_{\mathbf{W}}\mathcal{L}=\partial\mathcal{L}/\partial\mathbf{W}$. Then, the gradient of $i$-th neuron (i.e., column) of the weight matrix after masking is denoted as $\nabla_{\mathbf{W}_{:,i}}\tilde{\mathcal{L}}=$ $\gamma$ $\odot$ $\nabla_{\mathbf{W}_{:,i}}\mathcal{L}$, where $\gamma\in\{\mathbf{0}_{d},\mathbf{1}_{d}\}$ and $\odot$ means the Hadamard product. We also can mask bias terms similar to the weight matrix. Notice that this method is model-agnostic since all neural networks consist of linear transformation layers.
@@ 5.1 FaithUn Setups
Models. We adopt the instruction-tuned Gemma-2 (Gemma et al., 2024) models (2B & 9B) and the Llama-3.2 (Dubey et al., 2024) model (3B) to evaluate unlearning methods since they are among the latest open-source language models showing excellent performance.
Data. We sample 5% as the forget set and 10% as the retaining set from the Base QA dataset $\mathcal{C}$ since there are generally fewer samples to unlearn than to retain in real-world scenarios. More experiments on varying numbers of samples for the forget set are shown in Appendix B.5. We select 70% of $\mathcal{C}$ as the test set, guaranteeing it is completely separate from $\mathcal{C}_{f}$ and $\mathcal{C}_{r}$. For the MCQA evaluation (Section 3.4), we manually select the instruction and randomly sample two false answer options from possible answers for each relation $r$. The details of an example of the MCQA format and selecting false answer options are shown in Appendix B.1 and B.2, respectively. We also conduct experiments on various prompt templates, and the results are described in Appendix B.6.4.
Training Setups. When unlearning is applied to a language model, there is often a trade-off between unlearning knowledge (i.e., UA, UA, and MAf) and retaining the model’s overall knowledge (i.e., TA, SA, and MAt). Therefore, choosing the optimal model in the unlearning process is challenging since unlearning and retention are both important. For a fair comparison, we early stop the training procedure when UA0.33absent0.33$\leq$ 0.33≤ 0.33 is satisfied (random sampling from three options) to select the optimal model. More detailed experimental settings can be found in Appendix B.3.
5.2 Baselines
We adopt widely-used unlearning methods to evaluate the superficial unlearning: Gradient Ascent (GA), Gradient Ascent with a Retaining Loss (GAret), two Direct Preference Optimization variants (DPOmis and DPOrej), NPO (Zhang et al., 2024), and RMU (Li et al., 2024a). More details for the baselines are described in Appendix B.3. For KLUE, we select only 5% of neurons from Feed-forward networks for the knowledge neuron localization, and update them using general gradient ascent with retention loss. We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term. The experiments analyzing various hyper-parameters are shown in Section 5.5 and Appendix B.6.
@@ 5.2 Baselines
Table 2: Gemma-2 (2B) experimental results. We report the results of four metrics after unlearning the forget set (5%) in our settings. Bolded results indicate the best performance. We compute the accuracy over three trials and report the average accuracy.
  Case Method Questions for Forgetting Questions for Testing Label Logit Shift
@@
2 GAret "What is the country of citizenship of Ellen DeGeneres?" (Multi-hop QA) "What currency is associated with the country of citizenship United States dollar 0.5756 $\rightarrow$ 0.5757
@@
Table 3: Qualitative Analysis. GAret and KLUE are given the same questions to forget ($\mathcal{C}_{f}$) and test ($\mathcal{C}_{p}$, $\mathcal{C}_{m}$, and $\mathcal{C}_{s}$). Red texts indicate questions that should be forgotten, while blue texts should be retained. The "Label" and "Logit Shift" columns represent the golden labels for test questions and the logit changes corresponding to the labels, respectively. The underlined logit values indicate that the unlearning result is successful.

### v2

@@ 3.3 Dataset Summary
Dataset Statistics. After collecting triples of the Base QA dataset, we filter only triples including matched Multi-hop QA or Same-answer QA samples. Therefore, each QA instance in the Base QA dataset serves as a cluster for evaluating the faithfulness of unlearning methods. Consequently, we collect 664 QA pairs for the Base QA dataset. Each Base QA instance includes three paraphrased questions, for a total of 1,992 paraphrased QA instances in our dataset. FaithUn also include 1,714 instances for multi-hop QA datasets. Furthermore, our dataset includes 4,671 instances for the Same-answer QA dataset. The statistics of the constructed FaithUn datasets are shown in Table 7. We also describe detailed examples in Table 12.
Dataset Quality. We adopt a ChatGPT variant to generate natural language questions, a commonly used and powerful approach, following existing studies (Shi et al., 2024; Jin et al., 2024; Maini et al., 2024). However, to further investigate the quality of the dataset, we conducted a human evaluation of the generated questions. Specifically, we recruited crowd workers fluent in English through the university’s online community and had them evaluate 800 generated natural language questions. The results revealed an error rate of 0%, confirming the reliability of our benchmark.
@@ 3.4 Evaluation Framework
To evaluate the faithfulness of unlearning methods, we first split the forget set, the retain set, and the test set from the entire Base QA dataset. Then, we train LLMs to unlearn the forget set while maintaining knowledge of the retain set. We further evaluate the unlearned model on the test set to assess the knowledge for unseen data. In addition, we evaluate it on other datasets—the Paraphrased, Multi-hop, and Same-answer QA datasets—mapped to the forget and test sets.
Our unlearning framework consists of two types of input formats: (1) general QA format, and (2) multiple-choice QA (MCQA) format. We use the general QA format for unlearning and the MCQA format for evaluation. The general QA format inputs a question without an additional template, while the MCQA format uses a template that includes instructions and answer options. Suppose we aim to unlearn the knowledge of the question "Who is the mother of Barack Obama?", then we train an LLM not to output the correct answer (i.e., "Stanley Ann Dunham") using only the question as an input. However, many users use LLMs with various instruction templates, and an unlearned model should be evaluated in a stricter environment, considering generalization. Furthermore, assessing all possible answers to a question is one of the most challenging aspects of QA evaluation. Therefore, we utilize the MCQA form to assess an unlearned model. This makes it easier for LLMs to derive knowledge since they are given answer options; thus, it makes unlearning algorithms harder to apply. For this reason, we use the MCQA form to assess unlearned models in more challenging and practical settings. The details for the MCQA setting are described in Appendix B.1 and B.2.
@@ 3.5 Evaluation Metrics
We propose five metrics to evaluate the basic unlearning and the superficial unlearning performance. We use exact match to calculate the score of all metrics. (1) Unlearning Accuracy (UA): The accuracy for the forget set to evaluate the basic unlearning performance. (2) Extended Unlearning Accuracy (UA): The accuracy for the Paraphrased QA set to evaluate the generalized unlearning performance. (3) Test Accuracy (TA): The accuracy for the test set to evaluate whether knowledge of unseen data is maintained after the unlearning process. (4) Same-answer Test Accuracy (SA): The accuracy for the Same-answer QA set to analyze shortcut unlearning. An unlearning algorithm may only superficially degrade the probability of the answer regardless of context, as a trivial solution. (5) Multi-hop Test Accuracy (MA): The accuracy for the Multi-hop QA set matched with each instance of the forget set and test set to evaluate whether the interconnected knowledge of instances is effectively unlearned. We first derive individual accuracies for the multi-hop questions mapped to the forget set and test set, respectively. We denote MAf as the accuracy for the multi-hop questions mapped to the forget set, and MAt for the multi-hop questions mapped to the test set. Then, we compute the aggregated score, MA, by averaging the scores, (100100-MAf) and MAt. Although the number of samples in the test set is generally larger than in the forget set, we compute the average scores with equal weight, based on the assumption that unlearning the forget set is critical due to significant privacy concerns. (6) Total Score (Score): We average all the evaluation scores, (100100-UA), TA, SA, and MA, to present the overall performance.
@@ 4 Method: KLUE
In this section, we describe the method, KLUE, that identifies neurons contextually related to the target knowledge and updates only them during the unlearning process.
@@ 4.1.1 Knowledge Quantification
We utilize an attribution method (Yang et al., 2023) to extract the importance of neurons for specific world knowledge from LLMs. Formally, suppose we have $P_{\theta}(y|x)=\prod_{t=1}^{T}P_{\theta}(y_{t}|x,y_{1},...,y_{t-1})$ that represents a language model. The contribution of an $i$-th neuron to the representation $h$ in a particular layer, in predicting an answer $a$ given a question $q$ using $P_{\theta}$, is defined as follows:
@@
where $h^{l}$ means $l$-th token representation of $h$, and $\partial P_{\theta}(a|q)/\partial h^{l}_{i}$ is the gradient of $P_{\theta}(a|q)$ with respect to $h^{l}_{i}$. We use transformer variants for experiments; thus, activation scores and gradients of a specific layer are computed for each input token. Therefore, if an input text includes $L$ tokens, we have $L$ attribution scores for each neuron; thus, we aggregate attributions of tokens by using max pooling to acquire a single neuron attribution $A^{(q,a)}_{i}(h)$.
@@ 4.1.2 Superficial Knowledge Regularization
We identify one of the primary reasons that unlearning methods fail to operate faithfully as their tendency to adopt a trivial solution, namely shortcut unlearning, which reduces the probability of the target answer without considering the context. If the attribution score is computed solely based on the original question-answer pair $(q,a)\in\mathcal{C}$ targeted for unlearning, there is a potential risk that the method may select neurons that simply increase the likelihood of the answer $a$, regardless of context. To address this, we introduce a novel method, superficial knowledge regularization, which effectively excludes neurons associated solely with the answer but grounded in irrelevant contexts. Specifically, we first construct synthetic mismatched QA pairs $(q^{\prime},a)\in\mathcal{C}^{\prime}$, where $q^{\prime}$ is randomly sampled without regard to $a$, while the answer remains the same as the target answer. Then, we compute and average the attribution scores across all mismatched pairs. Consequently, we derive the final knowledge attribution $\mathcal{I}$, which captures only contextual knowledge by subtracting the mismatched attribution from the basic attribution, as follows:
$S^{(q,a)}_{i}(h)=\hskip-5.69046pt\sum_{(q^{\prime},a)\in\mathcal{C}^{{}^{\prime}}}{\hskip-7.11317pt\tilde{A}^{(q^{\prime},a)}_{i}(h)}$; (4)
@@
where $\mathcal{C}^{{}^{\prime}}$ is a set including mismatched question and answer pairs. $N$ is the number of mismatched samples, and $\alpha$ is a hyper-parameter to determine the magnitude of knowledge exclusion. $\tilde{A}$ means a negative value of $A$ is converted to the zero value. Since the negative values of the attribution are negative contributions to specific knowledge, we eliminate that unnecessary information. We use the forget and retain sets as a pool to sample mismatched questions. This approach mitigates the risk of shortcut unlearning, thereby naturally aligning with the goals of contextual unlearning. Notice that alleviating unlearning behaviors that disregard context inherently aligns with the objective of contextual unlearning; thus, it can improve multi-hop reasoning and mitigate shortcut unlearning.
4.2 Unforgotten Sample Unlearning
If we repeatedly unlearn samples that have already been sufficiently unlearned, it leads to overfitting in language models. Therefore, in each epoch’s unlearning process, we select and unlearn only questions that satisfy the knowledge memorization criteria (Described in Section 3.1).
4.3 Knowledge-localized Unlearning
After selecting unforgotten samples, we localize and update only the knowledge neurons corresponding to those selected samples in an LLM. Specifically, we first compute gradients of parameters for the selected unforgotten samples. Then, we quantify the knowledge relevance of each neuron by using the equations 3 and 4, and sort neurons of the whole target layers by the knowledge relevance scores; then, we select the top-$n$ knowledge neurons. We finally mask gradients of the parameters for knowledge-irrelevant neurons to exclude them from the unlearning process.
@@ 5.1 FaithUn Setups
Models. We adopt the instruction-tuned Gemma-2 (Gemma et al., 2024) models (2B & 9B) and the Llama-3.2 (Dubey et al., 2024) model (3B) to evaluate unlearning methods. These models serve as excellent starting points for unlearning evaluation, given their high default accuracy (above 80%) on the real-world entity QA benchmark.
Data. We sample 5% as the forget set and 10% as the retain set from the Base QA dataset since there are generally fewer samples to unlearn than to retain in real-world scenarios. We select 70% of $\mathcal{C}$ as the test set, guaranteeing it is completely separate from the forget and retain sets. For the MCQA evaluation (Section 3.4), we manually select the instruction and randomly sample two false answer options from the possible answers for each relation $r$. The details of an example of the MCQA format and selecting false answer options are shown in Appendix B.1 and B.2, respectively. We also conduct experiments on various prompt templates, described in Appendix C.4.
[…]
Baselines. We adopt widely-used unlearning methods to assess the superficial unlearning: Gradient Ascent (GA), Gradient Ascent with a Retention Loss (GAret), two Direct Preference Optimization variants (DPOmis and DPOrej), NPO (Zhang et al., 2024), and RMU (Li et al., 2024). Appendix B.3 describes more details for the baselines. For KLUE, we select only 5% of neurons from Feed-forward networks for the knowledge localization, and update them using general gradient ascent with retention loss. We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term. The experiments analyzing various hyper-parameters are shown in Section 5.4 and Appendix C.
Gemma-2
@@
Table 2: Gemma-2 (2B) results. We report the results after unlearning the forget set (5%) in our settings. Bolded results indicate the best performance. We report the average accuracy over three trials.
Llama-3.2
Default 90.91 87.28 85.65 50.57 -
GA 35.35 54.52 39.19 52.45 52.70
GAret 48.14 68.24 57.71 53.94 57.94
DPOrej 46.80 69.68 55.86 54.02 58.19
DPOmis 36.02 64.87 43.21 51.56 55.91
KLUE 45.79 77.58 65.12 53.99 62.73
Gemma-2
Default 91.92 89.87 86.57 48.07 -
GA 29.29 40.52 30.56 50.46 48.06
GAret 45.45 83.84 68.52 50.72 64.40
DPOrej 41.41 75.32 59.72 47.02 60.16
DPOmis 36.36 63.15 43.06 55.45 56.32
KLUE 40.40 89.83 81.48 60.48 72.85
Table 3: Llama-3.2 (3B) and Gemma-2 (9B) results. We report the results unlearning the forget set (5%).
1%
Default 72.22 85.34 71.43 54.18 -
GA 44.44 77.80 57.14 49.43 59.98
GAret 34.33 85.78 59.52 58.38 67.33
DPOrej 44.44 72.84 54.76 51.79 58.73
KLUE 36.11 85.34 63.09 59.77 68.02
10%
Default 83.84 85.34 76.82 50.05 -
GA 38.38 28.02 31.13 50.41 42.79
GAret 40.40 62.50 65.12 54.21 60.35
DPOrej 34.85 45.26 42.38 51.29 51.02
KLUE 40.91 81.03 69.98 59.18 67.32
Table 4: Gemma-2 (2B) results for varying forget sample sizes (i.e., 1% and 10%). The results for 5% is also shown in Table 2
  Case Method Questions for Forgetting Questions for Testing Label Prob Shift
@@
2 GAret "What is the country of citizenship of Ellen DeGeneres?" (Multi-hop QA)
"What currency is associated with the country of citizenship
United States dollar 0.5756 $\rightarrow$ 0.5757
@@ 2 GAret "What is the country of citizenship of Ellen DeGeneres?" (Multi-hop QA)
Table 5: Qualitative Analysis. GAret and KLUE are given the same questions for forgetting and testing. Red texts indicate questions that should be forgotten, while blue texts should be retained. The "Label" and "Prob Shift" columns represent the golden labels for test questions and the probability changes of the labels, respectively. and  indicate successful and failed unlearning, respectively.